# Coherently Coupled Carrier and Phonon Dynamics in Elemental Tellurium Probed by Few-Femtosecond Core-Level Transient Absorption

### v1

@@ Abstract
The narrow bandgap semiconductor elemental tellurium (Te) has a unique electronic structure due to strong spin-orbit splitting and a lack of inversion symmetry of it’s helical lattice. Using broadband extreme ultraviolet core-level transient absorption, we measure simultaneously the coherently coupled photo-induced carrier and lattice dynamics at the Te N4,5 edge initiated by a few-cycle NIR pulse. Ultrafast excitation of carriers leads to a coherently excited A1g phonon oscillation and the generation of a hot carrier population distribution that oscillates in temperature, and the phonon excursion and hot carrier temperature are out of phase with respect to each other. The depths of modulation suggest a significant coupling between the electronic and lattice degrees of freedom in Te. A long-lived shift of the absorption edge suggests a metastable excited state of Te in a new equilibrium potential energy surface that lives on the order of the carrier recombination timescale. The observed phonon-induced oscillations of the hot carriers are supportive of a semiconductor-to-metal light-induced phase transition, whereby Te becomes more metallic with increasing phonon-induced displacement. Additionally, near the Fermi level we observe an energy-dependent phase of the displacive excitation of the A1g phonon mode. The discovery of coherent coupling between the lattice and hot carriers in Te provides the basis to investigate coherent interactions between spin and orbital degrees of freedom. The results spectrally and temporally resolve the correlation between photo-excited hot carriers and coherent lattice excitations, providing insight on the optical manipulation of the tellurium electronic structure at high carrier densities exceeding .
@@ I Introduction
Tellurium (Te) is an elemental semiconductor that has attracted extensive interest due to its chiral structure of connected Van der Waals helical chains [1]. The strong spin-orbit interaction and chiral structure lead to many exotic electronic properties that include radial spin textures [2], topological phase transitions [3], and a quantum Hall effect [4]. An exceptionally high field effect mobility [5], resistance to oxidation [5], and thickness-tunable band gap of Te nanowires enables potential application in p-type transistors at the quantum confinement scale [6]. The high level of nested electronic bands also inherently leads to favorable thermoelectric properties in p-type Te where knowledge of carrier-phonon scattering is of vital importance [7]. Underlying these favorable properties are two anisotropic lone electron pairs at the valence band maximum (VBM) and isotropic anti-bonding -orbitals at the conduction band minimum (CBM). Important to realizing applications with Te is understanding the non-equilibrium carrier dynamics in both its valence and conduction bands. Manipulation of the chiral structure is possible by excitation of the electrons, which thus enables coupling of the electronic, lattice, and spin degrees of freedom. This is achievable via intense ultrafast laser pulses that may induce dynamic coupling between the electrons and lattice [8]. While optical-pump probe [9, 8, 10] and photoelectron spectroscopy [11] have made progress in the measurement of lattice and carrier dynamics in Te, simultaneous measurement remains difficult. In these regards, a time-resolved spectroscopic method that can resolve both the hot carriers and excited lattice configuration will be helpful in understanding the coupling of the lone pair and anti-bonding orbitals to obtain useful properties.
@@
In previous ultrafast studies of elemental Te, laser excitation was shown to induce a transient semiconductor-to-metal transition by distortion along the A1g phonon coordinate [8, 10, 12]. It was also demonstrated, using polarized optical transient reflectivity spectroscopy, that there is a band-splitting in the conduction band minimum that emerges at the H-valley and decays over 30 ps by relaxation of the carriers to the ground state [13]. More recently, it was determined with transient reflectivity in the mid-IR spectral region that the Auger recombination at the H-valley occurs over the same timescale [14]. Coherent dynamics of the A1g phonon have been studied extensively [15], investigating the impacts of carrier density, diffusion [9], and electronic softening [16] of the mode. Further investigation is needed as to how the A1g coherent phonon motion couples to the hot carriers that influence these effects and how hot carriers are possibly associated with a light-induced phase transition.
@@
In this work, we provide experimental characterization of the hot carrier dynamics and their coupling to the lattice via broadband extreme ultraviolet (XUV) transient absorption measurements at the Te N4,5 edge of Te thin films. The core-level transitions from the Te shell to the Fermi level probe the carriers in the valence band structure, changes to the lattice, and transient modification of the electronic structure with few-femtosecond resolution. We: (1) track the carrier thermalization and cooling processes, (2) characterize coherent coupling of the lattice and electronic degrees of freedom, and (3) measure the photo-induced lattice distortion coupled to the electronic structure. By exciting Te with a few-femtosecond laser pulse that spans visible to NIR wavelengths (500-1000 nm), at carrier densities around , we observe the generation of hot holes and electrons in the electronic band structure and coherent excitation of the A1g phonon mode. The high energy carriers rapidly thermalize through electron-electron (5 fs) and electron-phonon (0.15(1) ps) scattering to a hot Fermi-Dirac distribution and subsequently cool over the course of 1.60(2) ps. Due to coupling between electronic and lattice degrees of freedom, excitation of the A1g mode coherently modulates the temperature of the hot carriers with an observed nearly phase difference between the hot carriers the phonon-driven displacement. We attribute this to a variation of the electronic heat capacity as Te becomes more metallic when the structure is symmetrized by the A1g mode. This coherent anti-correlation between phonon excursion and carrier temperature, predicted by Giret et al. [17] and previously observed in the A7 semimetals [18, 19] and CDW materials [20], is also now observed in a degenerate semiconductor. Further, we measure a long-lived distortion of the lattice along the A1g phonon mode coordinate that decays over 31(2) ps. This is consistent with the hypothesis that carrier recombination times in Te [14] determine the lifetime of the metastable phase. Lastly, we resolve the spectral phase of the coherent phonon, which provides insight into the effects of hot carrier relaxation in the generation of the A1g phonon and identify a rad phase at the frequency attributed to weak excitation of the E1TO phonon.
II Experimental
@@
In the experiments we used polycrystalline Te films that are electron beam evaporated (Lebow) onto 30 nm thick silicon nitride windows (Norcada). The Te thickness was confirmed by atomic force scanning tip microscopy to be approximately 100 nm (see Fig. 6 in Appendix B). The trigonal form of Te is expected to form following crystallization of initially amorphous films over the course of a few minutes at room temperature [21]. This was confirmed by x-ray diffraction of the Te samples on silicon nitride (see Fig. 7 in Appendix B). To avoid sample evaporation, ablation, or laser pulse induced amorphization, careful consideration was taken to reduce laser induced heating during the experiment by chopping both the pump and probe to a repetition rate of 100 Hz. The XUV pump-probe experiments were performed at room temperature with a broadband attosecond XUV probe generated by focusing a near single cycle visible to near infrared laser pulse into argon gas for High Harmonic Generation (HHG) [22]. The same optical laser pulse is split before the HHG generation region in a Mach-Zender interferometer where it is delayed in time relative to the XUV probe from -100 to 3000 fs in steps of 6.6 fs. A measurement at shorter timescales used 0.66 fs steps from -30 to 90 fs along with parallel XUV transient absorption measurements in neon to obtain the zero-delay and the instrument response of 6.4(5) fs (see Appendix C). The carrier density induced by the pump is calculated by using the absorption coefficient and reflectance at 800 nm [23]. Further details on the instrument and XUV source are in Appendix A.
@@ III Results
In this section we first present the XUV transient absorption spectra at a carrier density exceeding the photo-induced phase transition threshold [12]. Subsequently, the spectral contributions are decomposed to disentangle the electronic and lattice degrees of freedom. We use both a simple model and then a bi-linear decomposition for analysis. Finally, the temporal response of the spectral contributions are characterized and fitted with a multi-temperature model.
@@ III.1 XUV Transient Absorption
In Fig. 1(a), the ground state core-absorption spectrum at the Te N4,5 edge in the range of 37 to 47 eV is shown. The increase in absorption, corresponds to the (42.3 eV) and (40.7 eV) semi-core electrons excited to the conduction band (CB) edge. The well resolved spin-orbit doublet with a 1.6 eV splitting and smooth absorption at higher energies is in agreement with synchrotron measurements of Te [24]. The absorption onset ( 41 eV) would correspond to transitions from to the unoccupied -orbitals at the H-point followed by higher lying transitions across the entire Brillouin zone (BZ). At energies above the peak there are small features that correspond to forbidden transitions to -bands well above the conduction band minimum (CBM) [25]. Due to the strong core-hole screening at the Te N4,5 edge, as previously reported [26], the absorption structure should well reflect a core-hole modified density of states in the conduction band. The lifetime broadening from the core-hole (on the order of 0.5 eV [25]), however, precludes our ability to resolve specific valleys in the absorption. The position of the valence band maximum in Te from the is previously reported to be 40.4 eV [27] and in p-type Te, the Fermi level is pinned near the valence band maximum (VBM) [28]. We expect our films to be naturally p-type, typical of evaporated Te [21].
@@
To track the dynamics, the change in the core-level absorption from the ground state at the Te N4,5 edge is detected by measuring the transmitted XUV with and without a time-delayed optical pump pulse. The XUV transient absorption spectra are shown in Fig. 1(b) at a carrier density of 1.5 . Typical of XUV transient absorption of thin films, a delay independent background before time zero that is due to static heating is subtracted. Following the arrival of the pump pulse at time zero, there is an increase in absorption starting at 39 eV, which then decreases around 41 eV. This pattern is replicated by the spin-orbit doublet and persists beyond our measured time delay. Within the first picosecond, the beating pattern centered around 40.5 eV and 42 eV resembles a shifting back and forth of the absorption edge at a frequency of 3.15 THz. The beating is also visible above 43 eV in the bands far from the Fermi level. The beating frequency of the absorption edge is consistent with the coherent excitation of the A1g phonon, but shifted from the equilibrium frequency of 3.6 THz [13]. Furthermore, at 40.5 eV, the A1g phonon beating exhibits a tilt in time with respect to energy, indicative of a non-flat phase distribution of the coherent motion versus photon energy.
@@
In addition to the lattice dynamics shifting the absorption edge, an increase in absorption is present below the VBM at 40.4 eV. This may be attributed to effects of state blocking/filling, resulting from the transitions of the core level to newly unoccupied states in the VB [26]. We therefore attribute this transient signal to the photo-generated hot holes in the VB. At the earliest timescale of the measurement, the increase in absorption would at most extend down to 38.2 eV, but will narrow in time as high energy (hot) holes relax (See Appendix D). Compared to the holes, which are spectrally isolated below the VBM of the , a decrease in absorption above the Fermi level from hot electrons is not as readily discerned due to spectral congestion between the lattice and carrier features. In the next section we focus on isolating the spectral contributions of the A1g phonon mode and the hot carriers.
III.2 Decomposing the Spectral Contributions of the A1g Phonon and Carriers
We first turn to the predominant spectral features in the transient absorption spectra at 150 fs, corresponding to the first oscillation maximum of the phonon. Well demonstrated in previous core-level TA measurements of elemental semiconductors [29] and transition metal dichacolgenides [26, 30], the excited state can be decomposed into individual contributions:
incorporating the phonon and band-gap renormalization induced absorption edge shift (), excited state inhomogeneous broadening of the core-level transition (), and carrier induced state blocking (filling) in the conduction (valence) band (), also called state removal/state opening. In the region of the Te N4,5 edge we show the decomposition into these components in Fig. 2. To retrieve the individual contributions, a global optimization is performed at 1 eV around the Fermi level with a basin-hopping algorithm to account for a high dimensionality minimization and the statistical errors retrieved via the Hessian matrix.
In this simplistic model the phonon and band-gap renormalization absorption edge shift is optimized to a 589(6) meV uniform energy shift with respect to the ground state XUV absorption. An optimized 5.1(6) % scaling factor is applied to the ground state XUV absorption to obtain agreement with the A. We believe the scaling factor is due to non-uniform shifting of the conduction band across the BZ following photoexcitation [29, 18]. The shift includes both changes to the conduction band energy and 4d core-level binding energies that we cannot independently discern. The inhomogeneous broadening of the excited state is achieved by convolving the static absorption with a Gaussian. As hot carrier thermalization is expected to occur mostly within the first 100 fs, we approximate the carrier induced state filling and blocking as a Fermi-Dirac distribution over parabolic bands with lifetime broadening. This is replicated for the spin-orbit split core-levels where there is strong spectral overlap between the holes of the and electrons of the . From this, we estimate a shared electronic temperature of 3180(50) K for both holes and electrons at the maximum of the phonon displacement and a hole Fermi level of 39.68(4) eV, about 0.8 eV below the VBM. Due to spectral congestion at the electron state blocking signal, the assumption of a shared electronic temperature is made to reduce the parametrization of the optimization.
While reproducing the change in absorption near the Fermi level, the model deviates from experiment due to an overestimate of the edge shift at high energies (above 42 eV). This we attribute to a non-uniform electron-phonon coupling over a large photon energy range [29, 18]. Although this approach provides precise quantities within the approximations, lack of detailed knowledge of complex band dynamics and strong spectral overlap of lattice and carrier features gives rise to significant uncertainty. Further the necessity to utilize a scaling factor in the rigid shift decomposition due to an unknown dependence of the shift on band dispersion and energy impedes successful interpretations with this decomposition method. We therefore analyze the complete TA spectra with another unbiased decomposition method and apply the previous methodology to validate the results.
@@ III.2 Decomposing the Spectral Contributions of the A1g Phonon and Carriers
To further evaluate our observations of the edge shift, broadening, and carriers, a singular value decomposition (SVD) is utilized to decompose the distinct transient contributions with unique spectral and kinetic profiles. As noticeable in Fig. 2, each contribution of shifting, broadening, and state filling/blocking has a unique derivative behavior in the spectral domain. The SVD approach enables isolation of the three basic contributions to the TA without presumptive modeling or restriction in the spectral and temporal domain. Further, it provides noise reduction and can isolate spectral contributions arising from strongly overlapping phonon and carrier signals of the spin-orbit doublet splitting. The SVD is implemented by adding the static absorption from Fig. 1(a) to the A at every time point, forcing the zero order basis vector to be the non-time-dependent ground state absorption.
@@
We find the entire transient signal is largely reproduced by three singular value components consistent with the above model. The spectra of the SVD are shown in Fig. 3(b,c,d) where the spectral SVD vectors are added back to the XUV ground state absorption for comparison. The largest SVD component (Fig. 3(b)) comprises 56% of the SVD decomposition and resembles a shift of the ground state absorption (Fig. 3(a)) to lower energies. There is an additional appearance of spectral features above the peak that cannot be reproduced by shifting the ground state XUV absorption. This SVD component we assign to the absorption edge shift due to coherent phonon, band-gap renormalization, and changes to the equilibrium position of Te along the A1g phonon coordinate from photoexcitation. From previous ultrafast work in Te at similar carrier densities, we expect an exceptionally large magnitude shift of the Te equilibrium position on the order of 26 pm [8, 9]. This is about ten times larger than what is reported in the semimetal Bismuth (2.6 pm) [18] and thus it is not surprising the shift of the Te absorption edge in the XUV is much larger compared to Bi [18]. The major contributions to the absorption edge shift therefore originate from changes to the binding energies of the core-levels from changes in bond length and shifting of the conduction band to lower energies.
@@
The second largest singular value (28% of the signal) shown in Fig. 3(c) readily resembles an unshifted XUV ground state absorption with a large increase in absorption below the valence band maximum and suppression of the peak relative to the . The increase in absorption below the VBM from the Te must be due to opening of states in the valence band, while the suppression of the peak comes from electrons populating the conduction band following excitation. The state filling/blocking contributions retrieved from the simple model agrees with the increase in absorption below the VBM in the second SVD vector as well as a decrease in the transition to the CB, respectively. Nevertheless, this singular value consistent with hot carriers incorporates both the holes and electrons. We find a small contribution of excited state broadening in the SVD decomposition (Fig. 3(d)) which agrees with the simple model. Both phonons and electronic effects could give rise to broadening [31], however, the relative contributions cannot be disentangled from the spectrum. While included for completeness in the SVD analysis, the excited state broadening is a minor effect that is not analyzed in more detail.
III.3 Temporal Dynamics of the A1g Phonon and Carriers

### v2

# Coherently Coupled Carrier and Phonon Dynamics in Elemental Tellurium Probed by XUV Transient Absorption

@@ Abstract
The narrow bandgap semiconductor elemental tellurium (Te) has a unique electronic structure due to strong spin-orbit splitting and a lack of inversion symmetry of it’s helical lattice. Using broadband extreme ultraviolet core-level transient absorption, we measure simultaneously the coherently coupled photo-induced carrier and lattice dynamics at the Te N4,5 edge initiated by a few-cycle NIR pulse. Ultrafast excitation of carriers leads to a coherently excited A1 phonon oscillation and the generation of a hot carrier population distribution that oscillates in temperature, and the phonon excursion and hot carrier temperature are out of phase with respect to each other. The depths of modulation suggest a significant coupling between the electronic and lattice degrees of freedom in Te. A long-lived shift of the absorption edge suggests an excited state of Te in a new equilibrium potential energy surface that lives on the order of the carrier recombination timescale. The observed phonon-induced oscillations of the hot carriers are supportive of a change in the metallicity, whereby Te becomes more metallic with increasing phonon-induced displacement. Additionally, near the Fermi level we observe an energy-dependent phase of the displacive excitation of the A1 phonon mode. The discovery of coherent coupling between the lattice and hot carriers in Te provides the basis to investigate coherent interactions between spin and orbital degrees of freedom. The results spectrally and temporally resolve the correlation between photo-excited hot carriers and coherent lattice excitations, providing insight on the optical manipulation of the Te electronic structure at high carrier densities exceeding .
@@ I Introduction
Tellurium (Te) is an elemental semiconductor that has attracted extensive interest due to its chiral structure of connected Van der Waals helical chains [1] (Fig. 1(a)). The strong spin-orbit interaction and chiral structure lead to many exotic electronic properties that include radial spin textures [2], topological phase transitions [3], and a quantum Hall effect [4]. An exceptionally high field effect mobility [5], resistance to oxidation [5], and thickness-tunable band gap of Te nanowires enables potential application in p-type transistors at the quantum confinement scale [6]. The high level of nested electronic bands also inherently leads to favorable thermoelectric properties in p-type Te where knowledge of carrier-phonon scattering is of vital importance [7]. Underlying these favorable properties are two anisotropic lone electron pairs at the valence band maximum (VBM) and isotropic anti-bonding -orbitals at the conduction band minimum (CBM). Important to realizing applications with Te is understanding the non-equilibrium carrier dynamics in both its valence and conduction bands. Manipulation of the chiral structure is possible by excitation of the electrons, which thus enables coupling of the electronic, lattice, and spin degrees of freedom. This is achievable via intense ultrafast laser pulses that may induce dynamic coupling between the electrons and lattice [8]. While optical-pump probe [9, 8, 10] and photoelectron spectroscopy [11] have made progress in the measurement of lattice and carrier dynamics in Te, simultaneous measurement remains difficult. In these regards, a time-resolved spectroscopic method that can resolve both the hot carriers and excited lattice configuration will be helpful in understanding the coupling of the lone pair and anti-bonding orbitals to obtain useful properties.
@@
In previous ultrafast studies of elemental Te, laser excitation was shown to induce a transient semiconductor-to-metal transition by distortion along the A1 phonon coordinate [8, 10, 12]. The distortion introduced by the A1 phonon motion is depicted Fig. 1(b,c). It was also demonstrated, using polarized optical transient reflectivity spectroscopy, that there is a band-splitting in the conduction band minimum that emerges at the H-valley and decays over 30 ps by relaxation of the carriers to the ground state [13]. More recently, it was determined with transient reflectivity in the mid-IR spectral region that the Auger recombination at the H-valley occurs over the same timescale [14]. Coherent dynamics of the A1 phonon have been studied extensively [15], investigating the impacts of carrier density, diffusion [9], and electronic softening [16] of the mode. Further investigation is needed as to how the A1 coherent phonon motion couples to the hot carriers that influence these effects, and what role they may have in the light-induced renormalization of the electronic structure. Additionally, discerning the energetic coupling between the hot carriers and coherent phonon is essential to clarifying the origin of the prominent coherent phonon motion in Te, where it is unclear as to whether the electronic temperature or carrier population is the driving force [15].
@@
In this work, we provide experimental characterization of the hot carrier dynamics and their coupling to the lattice via broadband extreme ultraviolet (XUV) transient absorption measurements at the Te N4,5 edge of Te thin films. The core-level transitions from the Te shell to the Fermi level probe the carriers in the valence band structure, changes to the lattice, and transient modification of the electronic structure with few-femtosecond resolution. An example of possible pump and probe transitions in our experiment are shown in Fig. 1(d) over a portion of the band structure. We: (1) track the carrier thermalization and cooling processes, (2) characterize coherent coupling of the lattice and electronic degrees of freedom, and (3) measure the photo-induced lattice distortion coupled to the electronic structure. By exciting Te with a few-femtosecond laser pulse that spans visible to NIR wavelengths (500-1000 nm), at carrier densities around , we observe the generation of hot holes and electrons in the electronic band structure and coherent excitation of the A1 phonon mode. The high energy carriers rapidly thermalize through electron-electron and electron-phonon (196(8) fs) scattering to a hot Fermi-Dirac distribution and subsequently cool over the course of 1.59(3) ps. Due to coupling between electronic and lattice degrees of freedom, excitation of the A1 mode coherently modulates the temperature of the hot carriers with an observed nearly phase difference between the hot carriers the phonon-driven displacement. We attribute this to a variation of the electronic heat capacity as Te becomes more metallic when the structure is symmetrized by the A1 mode. This coherent anti-correlation between phonon excursion and carrier temperature, predicted by Giret et al. [17] and previously observed in the A7 semimetals [18, 19] and CDW materials [20], is also now observed in a degenerate semiconductor. Further, we measure a long-lived distortion of the lattice along the A1 phonon mode coordinate that decays over 33(5) ps. This is consistent with the hypothesis that carrier recombination times in Te [14] determine the lifetime of the excited lattice configuration. Lastly, we resolve the spectral phase of the coherent phonon, which provides insight into the effects of hot carrier relaxation in the generation of the A1 phonon.
II Experimental Setup
@@
In the experiments we used polycrystalline Te films that are electron beam evaporated (Lebow) onto 30 nm thick silicon nitride windows (Norcada). The Te thickness was confirmed by atomic force scanning tip microscopy to be approximately 100 nm (see Fig. 6 in Appendix B). The trigonal form of Te is expected to form following crystallization of initially amorphous films over the course of a few minutes at room temperature [21]. This was confirmed by x-ray diffraction of the Te samples on silicon nitride (see Fig. 7 in Appendix B). To avoid sample evaporation, ablation, or laser pulse induced amorphization, careful consideration was taken to reduce laser induced heating during the experiment by chopping both the pump and probe to a repetition rate of 100 Hz. The XUV pump-probe experiments were performed at room temperature with a broadband attosecond XUV probe generated by focusing a near single cycle visible to near infrared laser pulse into krypton gas for High Harmonic Generation (HHG) [22]. The same optical laser pulse is split before the HHG generation region in a Mach-Zender interferometer where it is delayed in time relative to the XUV probe from -100 to 3000 fs in steps of 6.6 fs. A measurement at shorter timescales used 0.66 fs steps from -30 to 90 fs along with parallel XUV transient absorption measurements in neon to obtain the zero-delay and the instrument response of 6.4(5) fs (see Appendix C). The carrier density induced by the pump is calculated by using the linear absorption coefficient and reflectance at 800 nm [23]. In the sample, the pump has an estimated peak power density of and corresponding field strength of . Further details on the instrument and XUV source are in Appendix A.
@@ III Results
In this section we first present the XUV transient absorption spectra at an estimated 1.5 carrier density. Subsequently, the spectral contributions are decomposed to disentangle the electronic and lattice degrees of freedom. We use an analytical model and a bi-linear decomposition for analysis. Finally, the temporal response of the spectral contributions are characterized and fitted with a multi-temperature model.
@@ III.1 XUV Transient Absorption
In Fig. 1(e), the ground state core-absorption spectrum at the Te N4,5 edge in the range of 37 to 47 eV is shown. The increase in absorption, corresponds to the (42.3 eV) and (40.7 eV) semi-core electrons excited to the conduction band (CB) edge. The well resolved spin-orbit doublet with a 1.6 eV splitting and smooth absorption at higher energies is in agreement with synchrotron measurements of Te [25]. The absorption onset ( 41 eV) would correspond to transitions from to the unoccupied -orbitals at the H-point followed by higher lying transitions across the entire Brillouin zone (BZ). At energies above the peak there are small features that correspond to forbidden transitions to -bands well above the conduction band minimum (CBM) [26]. Due to the strong core-hole screening at the Te N4,5 edge, as previously reported [27], the absorption structure should well reflect a core-hole modified density of states (DOS) in the conduction band. The lifetime broadening from the core-hole (on the order of 0.5 eV [26]), however, precludes our ability to resolve specific valleys in the absorption. The position of the valence band maximum in Te from the is previously reported to be 40.4 eV [28] and in p-type Te, the Fermi level is pinned near the valence band maximum (VBM) [29]. We expect our films to be naturally p-type, typical of evaporated Te [21].
@@
To track the dynamics, the change in the core-level absorption from the ground state at the Te N4,5 edge is detected by measuring the transmitted XUV with and without a time-delayed optical pump pulse. The XUV transient absorption spectra are shown in Fig. 1(f) at a carrier density of 1.5 . Typical of XUV transient absorption of thin films, a delay independent background before time zero that is due to static heating is subtracted (see details in Appendix A). Following the arrival of the pump pulse at time zero, there is an increase in absorption starting at 39 eV, which then decreases around 41 eV. This pattern is replicated by the spin-orbit doublet and persists beyond our measured time delay. Within the first picosecond, the oscillation pattern centered around 40.5 eV and 42 eV resembles a shifting back and forth of the absorption edge at a frequency of 3.15 THz. The oscillation pattern is also visible above 43 eV in the bands far from the Fermi level. The oscillation frequency of the absorption edge is consistent with the coherent excitation of the A1 phonon, but shifted from the equilibrium frequency of 3.6 THz [13]. Furthermore, at 40.5 eV, the A1 phonon oscillation exhibits a tilt in time with respect to energy, indicative of a non-flat phase distribution of the coherent motion versus photon energy.
@@
In addition to the lattice dynamics shifting the absorption edge, an increase in absorption is present below the VBM at 40.4 eV. This may be attributed to effects of state blocking/filling, resulting from the transitions of the core level to newly unoccupied states in the VB [27]. We therefore attribute this transient signal to the photo-generated hot holes in the VB. At the earliest timescale of the measurement, the increase in absorption would at most extend down to 38.2 eV, but will narrow in time as high energy (hot) holes relax (See Appendix E). Compared to the holes, which are spectrally isolated below the VBM of the , a decrease in absorption above the Fermi level from hot electrons is not as readily discerned due to spectral congestion between the lattice and carrier features. In the next section we focus on isolating the spectral contributions of the A1 phonon mode and the hot carriers.
III.2 Decomposing the Spectral Contributions of the A1 Phonon and Carriers
@@
To evaluate our observations in the TA spectra, a singular value decomposition (SVD) is utilized to decompose the distinct transient contributions with unique spectral and kinetic profiles. To validate this approach, an analytical model is presented in the Appendix D where the TA can be decomposed into an energy shift of the absorption edge, inhomogeneous broadening, and state filling/blocking [30]. The SVD approach enables isolation of the basic contributions to the TA without presumptive modeling or restriction in the spectral and temporal domain. Further, it provides noise reduction and can isolate spectral contributions arising from strongly overlapping phonon and carrier signals of the spin-orbit doublet splitting. The SVD is implemented by adding the static absorption from Fig. 1(e) to the A at every time point, forcing the zero order basis vector to be the non-time-dependent ground state absorption.
@@
We find the entire transient signal is largely reproduced by three singular value components consistent with the analytical model. The spectra of the SVD are shown in Fig. 2(b,c,d) where the spectral SVD vectors are added back to the XUV ground state absorption for comparison. The largest SVD component (Fig. 2(b)) comprises 56% of the SVD decomposition and resembles a shift of the ground state absorption (Fig. 2(a)) to lower energies. There is an additional appearance of spectral features above the peak that cannot be reproduced by shifting the ground state XUV absorption. This SVD component we assign to the absorption edge shift due to the coherent phonon, band-gap renormalization, changes to 4d core-hole potential, and displacement of the equilibrium position of Te along the A1 phonon coordinate from photoexcitation. From previous ultrafast work in Te at similar carrier densities, we expect an exceptionally large magnitude shift of the Te equilibrium position on the order of 26 pm [8, 9]. This is about ten times larger than what is reported in the semimetal Bismuth (2.6 pm) [18]. Consistent with the anticipated large Te equilibrium position displacement, an absorption edge shift of 580(30) meV is estimated from the analytical model (see details in Appendix D). This value however should be interpreted cautiously due to approximations in deconvolving the spectrum. While many effects may shift the absorption, the most clear is the decreasing conduction band energy [8, 31]. Through the A1 phonon induced distortion, this decrease in conduction band energy will red shift the absorption by decreasing the transition energy from the 4d levels to the DOS. Further, the A1 phonon mode and lattice heating will modulate the bond lengths. This influences the 4d level ionic repulsion, thus changing the 4d binding energy. The exact dependence on the A1 mode requires detailed calculations of the core-level potential gradient, however, the lattice heating red shifts the absorption edge as indicated by the pre-time-zero delay heating (see Fig. 5(a)). Given the increase (decrease) in intra(inter)-chain bond distance upon increasing the chain radius, and distance scaling of the electrostatic repulsion, the core-level gradient is likely to counteract the decreasing conduction band energy with a blue shift. In addition to lattice dynamics, fast carrier screening of the 4d levels [32], changes in the valence to 4d level coulomb repulsion [33], or changes in electron localization may also shift the edge in the absence of lattice motion [34].
@@
The second largest singular value (28% of the signal) shown in Fig. 2(c) readily resembles an unshifted XUV ground state absorption with a large increase in absorption below the valence band maximum and suppression of the peak relative to the . The increase in absorption below the VBM from the Te must be due to opening of states in the valence band (holes), while the suppression of the peak comes from electrons populating the conduction band following excitation. The state filling/blocking contributions retrieved from the analytical model agrees with the increase in absorption below the VBM in the second SVD vector as well as a decrease in the transition to the CB, respectively. Nevertheless, this singular value consistent with hot carriers incorporates both the holes and electrons. We find a small contribution of excited state broadening in the SVD decomposition (Fig. 2(d)). Both phonons and electronic effects could give rise to broadening [33], however, the relative contributions cannot be disentangled from the spectrum.
III.3 Temporal Dynamics of the A1 Phonon and Carriers